# Ferrotoroidal order of the monolayer ice in martyite

### v1

@@ V.1 Simulation conditions
where the Lennard-Jones (LJ) size and energy parameters $\sigma_{ij}$ and $\epsilon_{ij}$ are given by the geometric averages. The LJ parameters and charges for the atoms in martyite are listed in Table SLABEL:tb:lj.
@@ V.2 Movie of simulated water molecules in martyite
The movie of molecular dynamics of water molecules in martyite is available online [http://]. The temperature is lowered from 300 K to 10 K in steps of 10 K. One second in the movie corresponds to 1.2 ps. In the movie, one can observe the transition at $T_{\text{c1}}\approx 190$ K from a state in which water molecules vibrate vigorously and rotate randomly at around honeycomb lattice sites to a state in which they form six-membered rings through hydrogen bonding, followed by the transition at $T_{\text{c2}}$ from a state in which the six-membered rings dynamically change their shapes to a state in which the six-membered rings are fixed in the distorted shapes.
@@ V.3 Comparison of MD simulations with the water models TIP4P/ICE and TIP4P/2005
Fig. S 10: Number of hydrogen bonds per water molecule in martyite as a function of temperature. The water models are TIP4P/2005 and TIP4P/ICE.
@@
Fig. S 11: Potential energies of the systems with ferrotoroidal ice and toroidal ice structures. The initial configuration of the ferrotoroidal ice was artificially constructed, while that of the toroidal ice was that which spontaneously formed during the cooling process of the MD simulation. The error bars are shown at each temperature. The inset shows the results from 300 K to 0 K.
@@
Fig. S 12: Ratio of the deformed hexamers as a function of temperature. The results using (a) TIP4P/ICE and (b) TIP4P/2005 models are shown.
V.3.1 Formation of six-membered rings
@@
Both of the water models give qualitatively the same results for the formation of six-membered rings by hydrogen bonding. Figure S10 shows the number $n_{\text{HB}}$ of hydrogen bonds per water molecule as a function of temperature. In the case of TIP4P/2005, $n_{\text{HB}}$, which is slightly smaller at 300 K, increases most rapidly in the range of 190 to 200 K, which is about 20 K higher than in the case of TIP4P/ICE.
V.3.2 Toroidal and ferrotoroidal structures
@@
When the temperature is lowered from 300 K to low temperatures, e.g. 10 K, ferrotoroidal order does not spontaneously appear. However, the potential energy for the ferrotoroidal structure (prepared by design and then equilibrated at low temperatures) is lower than for the toroidal structure, as shown in Figure S11. This is also the case for TIP4P/2005.
V.3.3 Distributions of the deformed six-membered rings (hexamers)
@@
Figure S12 shows the ratio of deformed hexamers in the ferrotoroidal structure as a function of temperature. In the case of TIP4P/ICE, all the hexamers have the same shape with one long bond at low temperatures. In the case of TIP4P/2005, however, 60% of the hexamers have the shape with two long bonds and 40% have the same shape with one long bond. Because of the coexisting hexamers with one- and two-long bonds, the octadecamer formation is hindered with the TIP4P/2005 model. Therefore, the shapes of the hexamers and their distributions are sensitive to the choice of the two water models.

### v2

@@ V.1 Simulation conditions
where the Lennard-Jones (LJ) size and energy parameters $\sigma_{ij}$ and $\epsilon_{ij}$ are given by the geometric averages. The LJ parameters and charges for the atoms in martyite are listed in Table S5.
@@ V.2 Movie of simulated water molecules in martyite
The movie of molecular dynamics of water molecules in martyite is available online. The temperature is lowered from 300 K to 10 K in steps of 10 K. One second in the movie corresponds to 1.2 ps. In the movie, one can observe the transition at $T_{\text{c1}}\approx 190$ K from a state in which water molecules vibrate vigorously and rotate randomly at around honeycomb lattice sites to a state in which they form six-membered rings through hydrogen bonding, followed by the transition at $T_{\text{c2}}$ from a state in which the six-membered rings dynamically change their shapes to a state in which the six-membered rings are fixed in the distorted shapes.
@@ V.3 Comparison of MD simulations with the water models TIP4P/ICE and TIP4P/2005
Fig. S 8: Number of hydrogen bonds per water molecule in martyite as a function of temperature. The water models are TIP4P/2005 and TIP4P/ICE.
@@
Fig. S 9: Potential energies of the systems with ferrotoroidal ice and toroidal ice structures. The initial configuration of the ferrotoroidal ice was artificially constructed, while that of the toroidal ice was that which spontaneously formed during the cooling process of the MD simulation. The error bars are shown at each temperature. The inset shows the results from 300 K to 0 K.
@@
Fig. S 10: Ratio of the deformed hexamers as a function of temperature. The results using (a) TIP4P/ICE and (b) TIP4P/2005 models are shown.
@@
Both of the water models give qualitatively the same results for the hexamer formation by hydrogen bonding. Figure S8 shows the number $n_{\text{HB}}$ of hydrogen bonds per water molecule as a function of temperature. In the case of TIP4P/2005, $n_{\text{HB}}$, which is slightly smaller at 300 K, increases most rapidly in the range of 190 to 200 K, which is about 20 K higher than in the case of TIP4P/ICE.
@@
When the temperature is lowered from 300 K to low temperatures, e.g. 10 K, the ferrotoroidal order of H2O does not spontaneously appear. However, the potential energy for the ferrotoroidal structure (prepared by design and then equilibrated at low temperatures) is lower than for the toroidal structure, as shown in Figure S9. This is also the case for TIP4P/2005.
@@
Figure S10 shows the ratio of deformed hexamers in the ferrotoroidal structure as a function of temperature. In the case of TIP4P/ICE, all the hexamers have the same shape with one long bond at low temperatures. In the case of TIP4P/2005, however, 60 % of the hexamers have the shape with two long bonds and 40 % have the same shape with one long bond. Because of the coexisting hexamers with one- and two-long bonds, the octadecamer formation is hindered with the TIP4P/2005 model. Therefore, the shapes of the hexamers and their distributions are sensitive to the choice of the water model.
V.4 Correspondence between superlattice peak intensity and hexamer number
We calculated the two-dimensional structure factor of O atoms of water in martyite by evaluating
$S(\mathbf{k})=1+\left<\frac{1}{N}\sum_{i=1}^{N}\sum_{j\neq i}^{N}\exp[-i\mathbf{k}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})]\right>,$
where $\mathbf{r}$ and $\mathbf{k}$ are the position and reciprocal space vectors in two dimensions. Figure 11 shows the (a) high- and (b) low-temperature $S(\mathbf{k})$ in two dimensions. At high temperatures, water molecules cannot form stable hydrogen bonds with each other (Fig. 4(d), 300 K), and therefore their average positions constitute the fundamental honeycomb lattice. $S(\mathbf{k})$ at 300 K in Fig. 11(a) shows only Bragg peaks. However, at sufficiently low temperatures, water molecules are so hydrogen-bonded with each other that they form nothing but hexamers (Fig. 10), and superlattice peaks are observed. Figure 11(b) shows the superlattice peaks in $S(\mathbf{k})$ at 10 K. Plotted in Fig. 11(c) are the simulated superlattice peak intensities and the ratio of hexamers, both as a function of $T$. It is clearly shown that the peak intensities and the hexamer ratio correspond to each other: as $T$ is decreased, both increase rapidly starting at 150 K and plateau at 100 K. Therefore, we conclude that the onset of the superlattice peaks results from the formation of hexamers of hydrogen-bonded water molecules.
Fig. S 11: (a, b) Two-dimensional structure factor at (a) 300 K and (b) 10 K. (c) The peak intensities at the indices of 4/3 1/3 0 (orange) and 5/3 2/3 0 (cyan) are plotted as a function of temperature. The black curve (right axis) shows the ratio of hexamers, including all types of deformed ones.
VI Critical temperatures
We comment on the slight discrepancy of $T_{\mathrm{c1}}$ observed in the dielectric (170–200 K) and XRD (200 K) experiments. The primary order parameter of the HT-LT phase transition is the electric toroidal dipole moment $\mathbf{A}$, which is not a polar multipole. Therefore, the dielectric measurement does not necessarily detect this ordering. The dielectric anomalies observed at $T_{\mathrm{c1}}$ (Fig. 2) probably reflect the ferrotoroidal domain dynamics, since the domain boundaries can locally possess electric dipole moments [44]. Similarly, the deformed hexamers below $T_{\mathrm{c2}}$ possess electric dipole moments, and the alignment of the type-1 hexamers near the domain boundaries can be influenced by external electric fields. The frequency dependences of these dielectric anomalies ($T_{\mathrm{c1}}$ and $T_{\mathrm{c2}}$) indicate the energy barrier to move these domain boundaries. Therefore, the thermodynamical transition temperatures are expected to be located at the beginning of the dielectric relaxations, namely, $T_{\mathrm{c1}}\approx 200$ K and $T_{\mathrm{c2}}\approx 50$ K.